OpenStudy (anonymous):

The vertices of the parallelogram are A(-1, -2), B(3, -2), C(4, 3) and D(0, 3). Find the area of the parallelogram. a. 16 units2 c. 30 units2 b. 20 units2 d. 3 units2

OpenStudy (mathmate):

A(-1, -2), B(3, -2), C(4, 3) and D(0, 3)  In the diagram, you see that CD is parallel to AB because they are both horizontal. C,D have y=3, A,B have y=-2, so height = 3-(-2)=5 Also, length of AB=(3-(-1))=4, and length of CD=(4-0)=4. With base (AB)=4 and height =5, the area can be calculated.